OpenStudy (angelmarie):

A right triangle can also be a(n) _____ triangle. scalene equilateral isosceles obtuse

OpenStudy (ajspeller):

This one has 2 answers X --- obtuse angles are greater than 90 X --- equilateral, no possible to have 3, 90 degree angles in a triangle

OpenStudy (danjs):

You can eliminate a couple to begin... Obtuse - a right triangle has 1 90 degree angle, and 2 other angles that add to 90. So it can not be obtuse. Equilateral - Pythgorean theorem a^2 + b^2 = c^2, cant have a=b=c